OpenStudy (anonymous):

Find the LCM of 8, 12, and 16. A. 48 B. 96 C. 192 D. 1536

OpenStudy (acxbox22):

which smallest number is divisible by all three numbers?

OpenStudy (anonymous):

\[78=2^3,12=2^2\times 3,16=2^4\] take each factor you see to the highest power you see it in any one number

OpenStudy (anonymous):

lol i meant \[8=2^3,12=2^2\times 3,16=2^4\]

OpenStudy (acxbox22):

:)

OpenStudy (acxbox22):

@satellite73 have you ever fanned anyone? :P

OpenStudy (anonymous):

take each factor you see to the highest power you see it in any one number is right however this is how you find the least common multiple of any numbers

OpenStudy (anonymous):

sally rand

OpenStudy (triciaal):